A student has an 80, 75, and 86 on 3 tests. What must her grade be on the next test if her average for the four tests is to be at least an 82?
Answer by checkley77(12844)

(80+75+86+x)/4=82
(241+x)/4=82
241+x=4*82
241+x=328
x=328-241
x=87 test score on ther last test to average 82%.